OpenStudy (dylandoty):

Which line is not perpendicular to the line 4x − 3y = 2? 3x + 4y = 2 3x − 4y = 2 -3x − 4y = 5 3x + 4y = -8

OpenStudy (aussiebabe23):

Get y by itself for all of the equations y=(2-3x)/4 once you do that for all of the equations, plug the first equation into a graphing calculator and then plug in each of the other graphs one at a time to compare. you are looking for the one that doesn't make some sort of T shape

OpenStudy (shrikantss9):

First get all equations to form y = ax + b , where a & b are numbers. Then the line perpendicular to the given line has value of 'a' such that it is negative of reciprocal to given equation's 'a' (i.e. -1/a) In this case , for equation 4x - 3y = 2, a = 4/3 therefore line perpendicular will have a = -3/4 equation 'a' 3x + 4y = 2 -3/4 3x - 4y = 2 3/4 -3x -4y = 5 -3/4 3x + 4y = -8 -3/4 as you can see, line 3x - 4y = 2 is not perpendicular to giiven line.